prove that cot(x/2)-3cot(3x/2) = ( 4sinx )/( 1+2cosx ).
 

Grade:11

1 Answers

Abhishek
6 years ago
3(cos^2/sin^2)-4/sin=1 3(cos^2/sin^2) - 4sin/sin^2 = 1 3cos^2 -4sin =sin^2 3cos^2-(1-cos^2) =4sin 4cos^2 -1 =4sin Cos^2 - sin=1/4 (1-sin^2) - sin =1/4 -Sin^2 - sin =-3/4 ...